OpenStudy (anonymous):

I need to chart my results of an experiment. ind variable is egg, and 1st dep variable is helment I made to protect it. 2nd dep variable is drop height, 12 and 24 inches. 1st 3 had type 1 helmet and dropped 12 in with 1 of 3 broken, 2nd 3 same helment dropped 24 inches all broken. Second type of helment 3 dropped 12 in no breaks, and same at 24 in how do I chart this? I am in 6th grade this is for my science fair project.

OpenStudy (compassionate):

Hi, hun, is it this one?

OpenStudy (anonymous):

yes thank you soooo much

OpenStudy (anonymous):

stepped on computer

OpenStudy (compassionate):

Independent = Egg Dependent = Helmet Dependent #2: Drop Height Helmet 1 #2 = (0, 12), (0 , 24) 12 inches = 1/3 broken 24 inches = 3/3 Helmet 2 12 inches = 0/3 24 inches = 0/3 ____________________ Can you do it from here, hun?

OpenStudy (anonymous):

so just plot it on the coordinate plane? I was not sure because bar graph , circle graph and scatter plot were also mentioned in the instructions on data at teh sciencebudedies website. but this makes sense. Thanks so much.

OpenStudy (compassionate):

Well, hun, you can also use a bar graph. Remember, the rates that remain the same will go on the bottom, and the changing ones go on the side.! :) I hoped this helped! Thank you for using OpenStudy, and if you EVER need help, feel free to contact me or the men in purple :)